# Using Biot-Savart to shrink Eulerian domains while maintaining or improving external flow accuracy

### v1

@@ 3 Multilevel reconstruction validation
We start by validating the multilevel reconstruction algorithm presented above. We selected to cases such respect our assumption of vorticity being confined far from the domain’s boundary. We demonstrate the 2D and 3D reconstruction error of the multilevel and the speedup generated by selecting the correct kernel radius from the above analysis.
Our Biot-Savart boundary conditions are implemented inside WaterLily.jl [^2] https://github.com/weymouth/WaterLily.jl, a fast, immersed-boundary flow solver based on the Boundary data immersion method [3] that can execute both on CPU and GPU [20]. To leverage the GPU capability of the flow solver and fully benefit from our novel boundary conditions, we perform all computations presented herein on an NVIDIA-A6000 GPU (48GB memory). All the codes used for this study and the instructions to reproduce these results are freely available at https://github.com/weymouth/BiotSavartBCs.jl.
3.1 2D validation: Lamb vortex
The velocity field generated by the Lamb-Chaplygin dipole is obtained from the scalar stream function $\vec{u}=-\vec{e}_{z}\times\nabla\psi$ given by
@@
where $J_{0}$ and $J_{1}$ are the zeroth and first Bessel functions of the first kind, respectively. $\beta$ is a parameter with the value $\beta=1.2197\pi/R$. The core of the vortex is defined by the isoline $\psi(r=R)=0$, that delimits the inside and outside of the vortex.
(a) $\log(\text{size})=1$
(b) $\log(\text{size})=3$
(c) $\log(\text{size})=5$
(d) $\log(\text{size})=8$
Figure 3: $L_{2}$-norm of the error in reconstructed velocity field outside the core of the 2D Lamb-Chaplygin vortex for varying (multi-level) kernel size. Error bars are shown in log-scale.
We reconstruct the velocity field outside the vortex core using the internal vorticity and our multilevel Biot-Savart method and compare that to the analytic result. We reconstruct the velocity on a domain of size $[2^{8},2^{8}]$. The dependency of the absolute error in the reconstructed velocity field on the radius of the multilevel refinement criterion are shown in Fig. 3. Setting the refine size $\tilde{S}$ to the domain size $\cal D$ eliminates the multi-level-error, but the result is not exact since the analytic vorticity distribution is sampled on a finite grid of size $\Delta x=0.05R$. This error depends primarily on the distance from the vortex, Fig. 3(d). Progressively reducing the refinement size $\tilde{S}$ introduces the multilevel error quantified in the previous section, which is only weakly dependent on the distance to the vortex.
This is further quantified in Fig. 4(a), which shows the maximum error in the reconstructed velocity field at a distance $d$ from the vortex from the Lamb-Chaplygin vortex. The size $\tilde{S}$ is shown to dominate the reconstruction error, with fairly uniform total error levels once $\tilde{S}\leq 2^{3}=8$. Fig. 4(a) reports the measured computational speed-up achieved by the multilevel method relative to setting $\tilde{S}=\cal D$, demonstrating up to 300$\times$ acceleration.
@@
with the corresponding errors and speed-up shown in Figure 5. Note that the analytic vorticity of the hill vortex is discontinuous, increasing the truncation error near the vortex compared to the smooth Lamp-vortex, a problem which would not be observed with any real three-dimensional viscous flow.
As argued in Sec. 2.3 the multilevel error decreases extremely quickly as $\tilde{S}$ is increased for three-dimensional flows. This is due to the scaling of the Biot-Savart kernel itself. The 3D case demonstrates a stronger sensitivity to the distance to vortex, as expected from the error estimate. Similarly, the speed-up is much faster in 3D because each doubling gathers 8 cells instead of 4, with a max observed speed-up of nearly 20,000$\times$ using the multilevel method.
@@ 4 Confined vorticity applications
We next validate the application of the Bio-Savart conditions to the solution of a set of impulsive flow problems. The flow is truly irrotational outside the computational domain, matching the implicit assumptions required to use the Biot-Savart equations to reconstruct the velocity on the boundary.
4.1 3D flow around an accelerating disk
The flow around a disk accelerated from rest in a quiescent flow is purely potential for $t\to 0^{+}$. As a result, an analytical expression for the added mass coefficient and the pressure field can be used to validate our novel potential flow boundary condition when injected in a moment step of the Navier-Stokes equations. We place a disk of radius $R$ in an initially quiescent flow. The disk is prescribed with a constant acceleration $a$ during one non-dimensional time $t^{*}=at/U$
@@
We integrate the Navier-Stokes equation during a single time-step from $t=0\to 0^{+}$ and compute the pressure force acting on the disk for various domain sizes. From potential flow theory, the added-mass coefficient for an impulsively started flow is
$Ca=\frac{F_{am}}{\rho D^{3}a}=\frac{F_{am}}{\rho D^{2}U^{2}}\left(\frac{1}{a^{*}}\right)=\frac{1}{3}$ (33)
@@
where $Ca$ is the classical added-mass coefficient for a circular plate. We then continue the simulation until $t^{*}=3$. Once the disk reaches its peak velocity, the Reynolds number is $Re=\frac{UR}{\nu}=1000$.
@@
We note here that the problem possesses two axes of symmetry that would allow only $1/4$ of the domain to be modelled. This can be easily accounted for in the Biot-Savart boundary conditions by adding the contributions of the images of each point. However, with our multilevel approach, they might lie on different levels, rendering the evaluation cumbersome. To avoid unnecessary complications, we model the full plate and do not use the symmetries of the problem.
@@
Figure 6-7 shows the vorticity and pressure field at three consecutive times during the cylinder’s motion. Fig. 6 show the results with the standard reflective boundary conditions (the entire computational domain is shown), and in Fig. 7, the flow has been obtained with the Biot-Savart boundary conditions. High blockage effects are apparent for the reflective boundary conditions, with stronger shear layers and a vortex center located further downstream than the Biot-Savart results. The pressure field is even more striking. With iso-contour of the pressure sticking to the wall to satisfy the homogeneous Neumann condition set by the velocity field, see . Because the Biot-Savart boundary conditions improve this boundary condition, the pressure follows the expected potential flow solution and, more importantly, is relaxed from the homogeneous Neumann condition that reflective boundary conditions imposed.
Figure 6: Flow around an initially stationary disk accelerated in a quiescent flow at three convective times, $t^{*}\in[1,2,3]$, using the reflective boundary conditions. The entire computational domain is shown. The top half shows 10 iso-contour of the vorticity, equally spaced in the interval $\omega L/U\in\pm 20$. The bottom half shows the pressure field, with 10 iso-contour equally spaced in $p/\rho U^{2}\in\pm 2$.
Fig. 8 shows the time trace of the pressure force acting on the accelerated disk. The difference in the pressure field is found again in the difference in the forces. Here, the Biot-Savart boundary condition can exactly recover the added-mass force on the first time step, while the reflective boundary condition overestimates it. For larger $t^{*}$, the effect of the boundary condition is still observable; the reflective boundary condition significantly overestimates the pressure forces
@@
Figure 7: Flow around an initially stationary disk accelerated in a quiescent flow at three convective times, $t^{*}\in[1,2,3]$, using the Biot-Savart boundary conditions. The entire computational domain is shown. The top half shows 10 iso-contour of the vorticity, equally spaced in the interval $\omega L/U\in\pm 20$. The bottom half shows the pressure field, with 10 iso-contour equally spaced in $p/\rho U^{2}\in\pm 2$.
@@
Figure 8: Instantaneous pressure force acting on the disk for the two different boundary conditions. The convective time corresponds to the snapshots shown in Fig. 6-7.
@@
Figure 9 shows the vorticity generated by this accelerating body as six convective times $t^{*}\in[1,2,3,4,5,6]$. The vortices generated are shown via the $\lambda_{2}$-criterion [21]. The initial impulse generates a strong shear layer around the plate’s edge that induces a strong vortex. Although the platform of the plate is square, this initial vortex rearrange itself into a shape resembling that of a vortex ring. After this initial vortex has traveled downstream, the plate’s corner starts shedding smaller vortices that later trigger the breakdown of the start vortex. An animation of these still frames is provided in the supplementary material.
@@ 5 Vortex-street applications
Numerous examples of external flow violate our initial assumption that the vorticity is confined near the immersed body, far from the domain’s boundary. We use two classical examples of vortex-street flows to investigate the effect of our short computational domain on flow history. First, the flow around a 2D cylinder at a Reynolds number $Re=200$, which results in the classical van Kàrmàn wake. Then, the reversed van Kàrmàn (or propulsive) wake generated by a heaving airfoil in a uniform flow at a range of Strouhal numbers.
5.1 Flow around a 2D cylinder at $Re=200$
We start our analysis with the flow behind a 2D circular cylinder of diameter $D$. The cylinder is immersed in a viscous fluid with free-stream velocity $U=(1,0)$. The viscosity is set such that the Reynolds number is $Re=\frac{UD}{\nu}=200$. We perform a domain study by varying the domain sizes from $5D\times 3D$ up to $30D\times 24D$ and, for each domain, compare the reflective and Biot-Savart boundary conditions.
Figure 10 shows the total drag force acting on the cylinder for the various domain sizes and boundary conditions. The different domain sizes and the Biot-Savart boundary conditions yield the best results with an excellent collapse of the drag forces, regardless of the domain size. Simulations performed with the Biot-Savart boundary conditions are insensitive to domain size. Reflective boundary conditions induce extremely large blockage effects for small domains and only slowly converge to the Biot-Savart boundary solution for significantly larger domains.
Additionaly, we compare the results obtained with a grid aligned with the mean flow direction to results obtained on a $7L\times 7L$ domain where the flow direction is $\vec{U}=(1/\sqrt{2},1/\sqrt{2})$ with the Biot-Savart boundary conditions. Results are also presented in Fig. 10 and are almost indistinguishable from the results using the aligned domain.
In addition to the total drag forces acting on the cylinder, we look at the vortex shedding frequency in the wake. Here again, the Biot-Savart boundary conditions show an excellent collapse of the shedding frequency, which matches that of classical bluff body vortex shedding frequency of $St\sim 0.3$ The high blockage effect induced by the reflective boundary conditions induces fast flow around the cylinder, increasing the shedding frequency to higher values with reflective boundary conditions.
Figure 10: (left) Drag force acting on the circular cylinder at $Re=200$ for various combinations of boundary conditions and domain size. The rotated domain consists of a $4D\times 4D$ domain, with the cylinder centered at $1.5D\times 1.5D$ and a free stream velocity of $\vec{U}=(1/\sqrt{2},1/\sqrt{2})$. (right) The frequency spectrum of the wake of the cylinder for the same combination of boundary conditions and domain size shows the classical shedding frequency at $St=0.3$.
These results suggest that, for drag wakes at least, the wake history on the immersed body is very small. However, propulsive systems use thrust wakes where wake history is expected to be higher.
@@ 5.2 Wake behind a heaving foil
We continue our analysis with thrust wakes. We immerse a rigid foil of length $L$ moving with a linear velocity $\vec{U}=(U,0)$ in a viscous fluid. In addition to its steady forward speed, the foil undergoes a pure heave motion of amplitude $h_{0}$
@@
where $h_{0}/L=0.5$ is the non-dimensional amplitude, and the frequency is $f=St\,U/h_{0}$. For Strouhal number $St<0.5$, the airfoil generates thrust, and almost no net lift, and for $St>0.6$, the wake generated by this flapping foil is strongly asymmetric and generates a significant net lift force. We perform simulations of this system for a range of Strouhal numbers and compare the mean lift force and wake obtained on a large domain ($30L\times 20L$) with standard reflective boundary conditions to the ones obtained on two different domains of size ($6L\times 4L$) and ($12L\times 8L$) that use the Biot-Savart boundary conditions. We simulate the flow around this heaving airfoil for 100 convective times to ensure that the wake stabilizes, and we gather pressure and viscous forces for the last 20 convective times. Mean lift forces are obtained by time-averaging these forces over the past 20 convective times.
For $St<0.5$, we observe almost no difference between the different boundary conditions and domain sizes. The wake is symmetric, and the airfoil produces zero mean lift forces; see Fig. 12. For $St\geq 0.5$ we start to observe wake deflection on the large domains with the reflective boundary condition.
@@
Figure 11: Snapshot of the deflected wake behind an airfoil at $St=0.6$ and $Re=100$ for an amplitude to chord ratio $h_{0}/L=1.0$. The upper left panel $(a)$ shows the far wake obtained with the reflective boundary condition in a large domain, the upper right panel $(b)$ shows a zoom onto the airfoil and the near wake, the bottom left $(c)$ panel shows the entire (small) Biot-Savart domain used for the computation (the dashed line represents the domain boundaries) and the lower right panel $(d)$ shows the entire (larger) Biot-Savart domain. Vorticity is shown as 10 equally spaced isocontours in the interval $\omega L/U\pm 10$.
Fig. 11 shows the vorticity field generated by the wing at $St=0.6$. The left panel shows the complete wake obtained with the large domain of size ($30L\times 20L$). The left panel is a zoomed-in version of the same simulation, using a field of view with the size of the largest Biot-Savart domain. Finally, the bottom row shows two Biot-Savart simulations with two different domain sizes. The left panel shows the smallest domain of size ($6L\times 4L$), represented by the gray dashed line, and the right panel shows a larger domain of size ($12L\times 8L$) which extends to the whole panel.
Wake history strongly influences the near wake; the large domain correctly captures this, and a strong wake deflection is observed (the first vortex shed sets the direction; in this case, it is shed in the negative $y$ direction).
Figure 12: (left) Mean lift force acting on the airfoil at $Re=100$, for a range of Strouhal numbers for the Biot-Savart and the reflective boundary condition and (right) Poincaré map of the lift and drag coefficients for $St=0.6$
@@
While this is a limitation of our method, at least in terms of domain size, this flow is an artifact of 2D simulations and does not occur in actual flows. Indeed, experiments on finite foils have shown that these high-lift deflected jets cannot form for finite wings, see [22, 23].

### v2

@@ 3 Multilevel reconstruction validation
We validate the multilevel reconstruction algorithm presented above using 2D and 3D analytic vorticity fields. We also study the multilevel reconstruction error and the speedup generated by selecting the correct kernel radius from the above analysis.
Our Biot-Savart boundary conditions are implemented inside WaterLily.jl [^2] https://github.com/weymouth/WaterLily.jl, a fast, immersed-boundary flow solver based on the Boundary data immersion method [22] that can execute both on CPU and GPU [28]. To leverage the GPU capability of the flow solver and fully benefit from our novel boundary conditions, we perform all computations presented herein on an NVIDIA-A6000 GPU (48GB memory). The complete code and instructions needed to reproduce these results are freely available at https://github.com/weymouth/BiotSavartBCs.jl.
3.1 2D validation: Lamb Chaplygin vortex-dipole
The velocity field generated by the Lamb-Chaplygin dipole is defined by the scalar stream function
@@
where $J_{0}$ and $J_{1}$ are the zeroth and first Bessel functions of the first kind, respectively, and $\beta$ is a parameter with the value $\beta=1.2197\pi/R$. The 2D velocity field is given by $\vec{u}=-\vec{e}_{z}\times\nabla\psi$, and the core of the vortex is defined by the isoline $\psi(r=R)=0$.
$10^{-1}$
$10^{-2}$
$10^{-3}$
$10^{-4}$
$10^{-5}$
$10^{-6}$
$a)$ $S=2$
$c)$ $S=8$
$d)$ $S=16$
$f)$ $S=128$
$b)$ $S=4$
$e)$ $S=32$
Figure 3: $L_{2}$-norm of the error in reconstructed velocity field outside the core of the 2D Lamb-Chaplygin vortex for varying (multi-level) kernel size.
We reconstruct the velocity field outside the vortex core using the internal vorticity and our multilevel Biot-Savart method and compare that to the analytic result. We reconstruct the velocity on a domain of size $[2^{8},2^{8}]$. The dependency of the absolute error in the reconstructed velocity field on the radius of the multilevel refinement criterion are shown in Fig. 3. Setting the refine size $S$ to the domain size $\cal D$ eliminates the multi-level-error, but the result is not exact since the analytic vorticity distribution is sampled on a finite grid of size $\Delta x=0.05R$. This error depends primarily on the distance from the vortex, Fig. 3(f). Progressively reducing the refinement size $S$ introduces the multilevel error quantified in the previous section, which is only weakly dependent on the distance to the vortex.
This is further quantified in Fig. 4(a), which shows the maximum error in the reconstructed velocity field at a distance $d$ from the vortex from the Lamb-Chaplygin vortex. The size $S$ is shown to dominate the reconstruction error, with fairly uniform total error levels once $S\leq 2^{3}=8$. Fig. 4(b) reports the measured computational speed-up achieved by the multilevel method relative to setting $S=\Omega$, demonstrating up to 300$\times$ acceleration.
@@
with the corresponding errors and speed-up shown in Figure 5. Note that the analytic vorticity of the hill vortex is discontinuous, increasing the truncation error near the vortex compared to the smooth 2D Lamb-dipole.
As argued in Sec. 2.3 the multilevel error decreases extremely quickly as $S$ is increased for three-dimensional flows. This is due to the scaling of the Biot-Savart kernel itself. The 3D case demonstrates a stronger sensitivity to the distance to vortex, as expected from the error estimate. Similarly, the speed-up is much faster in 3D because each doubling gathers 8 cells instead of 4, with a max observed speed-up of nearly 20,000$\times$ using the multilevel method.
@@ 4 Confined vorticity applications
Next, we validate the application of the Biot-Savart conditions to a set of impulsive flow problems. The flow is truly irrotational outside the computational domain, matching the implicit assumptions required to use the Biot-Savart equations to reconstruct the velocity on the boundary.
4.1 2D and 3D flow around an accelerating disk
The flow around a 2D or 3D disk accelerated from rest in a quiescent flow is purely potential for $t\to 0^{+}$. As a result, an analytical expression for the added mass coefficient can be used to validate our novel potential flow boundary condition when injected in a moment step of the Navier-Stokes equations.
The disk of radius $R$ and diameter $D=2R$ is given a constant acceleration $a$ for the initial non-dimensional time $t^{*}=at/U\leq 1$ and is given a constant speed after this, ie
@@
We perform 2D and 3D simulations of the flow around a plate/circular disk immersed in a domain of dimension $1.5D\times 1.5D(\times 1.5D)$. From potential flow theory, the added-mass coefficient for an impulsively started flow is
$Ca=\frac{F_{am}}{\rho D^{3}a}=\frac{F_{am}}{\rho D^{2}U^{2}}\left(\frac{1}{a^{*}}\right)$ (26)
@@
where $Ca$ is the classical added-mass coefficient. We simulate the 2D or 3D flow until $t^{*}=3$. Once the disk reaches its peak velocity, the Reynolds number is $\text{Re}=\frac{UR}{\nu}=1000$.
@@
We note here that the 2D and 3D problems possess axes of symmetry that would allow only a portion of the domain to be modeled. This can be easily accounted for in the Biot-Savart boundary conditions by adding the contributions of the images of each vortex cell, see A. However, to avoid unnecessary complications in this manuscript, we model the full plate and do not use the problem’s symmetries.
@@
Figure 6 shows the vorticity and pressure field at three consecutive times during the disk’s motion on the midplane of the 3D circular disk. The top row shows the results with the standard reflective boundary conditions (the entire computational domain is shown), and the bottom row shows the flow obtained with the Biot-Savart boundary conditions. High blockage effects are apparent for the reflective boundary conditions, which generate a stronger shear layer that feeds a stronger vortex located further downstream than the Biot-Savart results. The error in the pressure field due to the homogeneous Neumann condition placed close to the immersed body generates pressure peaks twice as large as with the Biot-Savart boundary conditions.
@@
Figure 6: Flow around an initially stationary disk accelerated in a quiescent flow at three convective times, $t^{*}\in[1,2,3]$, using the reflective boundary conditions (top) and the Biot-Savart boundary condition (bottom). A slice through the middle of the entire computational domain is shown. The top half shows 10 iso-contour of the vorticity, equally spaced in the interval $\omega L/U\in\pm 20$. The bottom half shows the pressure field, with 10 iso-contour equally spaced in $p/\rho U^{2}\in\pm 2$.
Fig. 7 shows the time trace of the pressure force acting on the accelerated disk, as well as the equivalent 2D simulation on an accelerating flat plate for comparison. The difference in the pressure field is found again in the difference in the forces. Here, the Biot-Savart boundary condition can exactly recover the added-mass force on the first time step, while the reflective boundary condition overestimates it. For larger $t^{*}$, the effect of the boundary condition is still observable; the reflective boundary condition significantly overestimates the pressure forces for all $t$.
@@
Figure 7: Instantaneous pressure force acting on the disk for the two different boundary conditions. The convective time corresponds to the snapshots shown in Fig. 6
@@
We qualitatively compare the shape of the vortex ring forming behind our accelerating square plate with the flow visualization of the same cases in [29], who used dye coloring to visualize the wake forming behind the accelerating plate. Results are presented in Figure 9-8 for four instants during and after the acceleration phase.
$a)\,\,t^{*}=2.0$
$b)\,\,t^{*}=2.5$
$c)\,\,t^{*}=3.0$
$d)\,\,t^{*}=4.0$
@@
An excellent agreement between the simulations and the experiments is found for this qualitative verification, both in the oblique and downstream view from the plate. The initial vortex roll-up starts to break down due to the high-strength vortices generated by the plate’s edges. The initial roll-up is extremely similar between both visualization techniques, and the shape of the vortex ring is extremely well captured by the simulations despite the minimal domain size.
$a)\,\,t^{*}=2.0$
$b)\,\,t^{*}=2.5$
$c)\,\,t^{*}=3.0$
$d)\,\,t^{*}=4.0$
[…]
@@ 5 Vortex-street applications
Next, we study the accuracy of the new Biot-Savart conditions on flow cases with vortex wakes that are truncated by the domain boundary, introducing an additional modelling error. We use two classical examples of vortex-street flows: (1) the flow around a 2D cylinder at a Reynolds number $\text{Re}=200$, which results in the classical van Kàrmàn wake, and (2) the reversed van Kàrmàn (or propulsive) wake generated by a heaving airfoil in a uniform flow at a range of Strouhal numbers.
5.1 Flow around a 2D cylinder at $\text{Re}=200$
We start our analysis with the flow behind a 2D circular cylinder of diameter $D$. The cylinder is immersed in a viscous fluid with free-stream velocity $U=(1,0)$. The viscosity is set such that the Reynolds number is $\text{Re}=\frac{UD}{\nu}=200$. We perform a domain study by varying the domain sizes from $5D\times 3D$ up to $30D\times 24D$ and, for each domain, compare the reflective and Biot-Savart boundary conditions.
Figure 10 shows the total drag force acting on the cylinder for the various domain sizes and boundary conditions. The Biot-Savart boundary conditions yield the best results with an excellent collapse of the drag forcesregardless of the domain size, including the tiny $5Dx3D$ domain which captures no more than two wake vortices at any time. Reflective boundary conditions induce extremely large blockage effects for small domains and only slowly converge to the Biot-Savart boundary solution with a domain 48x smaller.
Additionally, we compare the results obtained with a grid aligned with the mean flow direction to results obtained on a $7D\times 7D$ domain where the flow direction is $\vec{U}=(1/\sqrt{2},1/\sqrt{2})$ with the Biot-Savart boundary conditions. Results are also presented in Fig. 10 and are almost indistinguishable from the results using the aligned domain.
In addition to the total drag forces acting on the cylinder, we look at the vortex shedding frequency in the wake. Here again, the Biot-Savart boundary conditions show an excellent collapse of the shedding frequency, which matches that of classical bluff body vortex shedding frequency of $St\sim 0.3$. The high blockage effect induced by the reflective boundary conditions induces fast flow around the cylinder, adding a non-physical domain-size dependance on the shedding frequency.
Figure 10: (left) Drag force acting on the circular cylinder at $\text{Re}=200$ for various combinations of boundary conditions and domain size. The rotated domain consists of a $4D\times 4D$ domain, with the cylinder centered at $1.5D\times 1.5D$ and a free stream velocity of $\vec{U}=(1/\sqrt{2},1/\sqrt{2})$. (right) The frequency spectrum of the wake of the cylinder for the same combination of boundary conditions and domain size shows the classical shedding frequency at $St=0.3$.
@@ 5.2 Wake behind a heaving foil
The previous results suggest that the wake history on the immersed body is very small for drag wakes. We next continue our analysis with thrust wakes, where the influence of wake history is expected to be stronger.
We immerse a rigid foil of length $L$ moving with a linear velocity $\vec{U}=(U,0)$ in a viscous fluid. In addition to its steady forward speed, the foil undergoes a pure heave motion of amplitude $h_{0}$
@@
where $h_{0}/L=0.5$ is the non-dimensional amplitude, and the frequency is $f=\text{St}\,U/h_{0}$. For Strouhal number $\text{St}<0.5$, the airfoil generates thrust, and almost no net lift, and for $\text{St}>0.6$, the wake generated by this flapping foil is strongly asymmetric and generates a significant net lift force. We perform simulations of this system for a range of Strouhal numbers and compare the mean lift force and wake obtained on a large domain ($30L\times 20L$) with standard reflective boundary conditions to the ones obtained on two different domains of size ($6L\times 4L$) and ($12L\times 8L$) that use the Biot-Savart boundary conditions. We simulate the flow around this heaving airfoil for 100 convective times to ensure that the wake stabilizes, and we gather pressure and viscous forces for the last 20 convective times. Mean lift forces are obtained by time-averaging these forces over the past 20 convective times.
@@
Figure 11: Snapshot of the deflected wake behind an airfoil at $\text{St}=0.6$ and $\text{Re}=100$ for an amplitude to chord ratio $h_{0}/L=1.0$. The upper left panel $(a)$ shows the far wake obtained with the reflective boundary condition in a large domain, the upper right panel $(b)$ shows a zoom onto the airfoil and the near wake, the bottom left $(c)$ panel shows the entire (small) Biot-Savart domain used for the computation and the lower right panel $(d)$ shows the entire (larger) Biot-Savart domain. Vorticity is shown as 10 equally spaced isocontours in the interval $\omega L/U\pm 10$.
Figure 12: (left) Mean lift force acting on the airfoil at $\text{Re}=100$, for a range of Strouhal numbers for the Biot-Savart and the reflective boundary condition and (right) Poincaré map of the lift and drag coefficients for $St=0.6$
@@
We expect that in any case where wake history effects are highly dominant, the new Biot-Savart boundary condition will require a domain large enough for this wake to develop “stable” dynamics before exiting the domain. However, we should note that the case of 2D deflected thrust wakes is somewhat pathological, depending on long-range vortex interactions only possible in 2D flows. Indeed, experiments on finite foils have shown that these high-lift deflected jets cannot form for finite wings, see [30, 31].